OpenStudy (anonymous):

Which of the following prefixes represents the smallest value? (2 points) Question 5 options: 1) Centi 2) Deca 3) Deci 4) Milli

OpenStudy (anonymous):

@mathstudent55

OpenStudy (mathstudent55):

centi = 1/100 deca = 10 deci = 1/10 milli = 1/1000

OpenStudy (anonymous):

milli @mathstudent55

OpenStudy (mathstudent55):

Correct.

OpenStudy (anonymous):

thank you @mathstudent55